Zia Co. makes flowerpots from recycled plastic in two departments, molding and packaging. At the beginning of the month, the molding department has 3,000 units in inventory, 70% complete as to materials. During the month, the molding department started 23,000 units. At the end of the month, the molding department had 4,500 units in ending inventory, 80% complete as to materials. Units completed in the molding department are transferred into the packaging department.
Cost information for the molding department for the month follows:

Beginning work in process inventory (direct materials) $ 2,200
Direct materials added during the month 37,900

QS 16-19 Weighted average: Assigning costs to output LO C3

Using the weighted-average method, assign direct materials costs to the molding department’s output—specifically, the units transferred out to the packaging department and the units that remain in process in the molding department at month-end. (Do not round intermediate calculations.)
Answer #1

Solution:

Nos of equivalent units of direct materials = Units completed and transferred out + Units in ending WIP * %of completion

= 21500 + 4500 *80% = 25100 units

Cost per equivalent unit = (Beginning WIP + Cost added during the period) / Nos of equivalent units = ($2,200 + $37,900) / 25100 = $1.59761 per unit

Cost of units transferred out to packaging department = 21500 * $1.59761 = $34,349

Cost of ending WIP = 3600*$1.59761 = $5,751